# UrbanGraph: Physics-Informed Spatio-Temporal Dynamic Heterogeneous Graphs for Urban Microclimate Prediction

### v1

@@ 5 Experiments
Dataset. All experiments are conducted on the high-fidelity dataset generated via ENVI-met, as detailed in Appendix B. Our primary task is to evaluate the model’s effectiveness on the UTCI prediction task. To further demonstrate the scalability and generalization capability of our physics-informed representation and architecture, we train and evaluate models on all six target variables. The temporal graph structures are constructed following the physical principles outlined in Section 4.1. The resulting spatio-temporal graph sequence data is split into training (70%), validation (20%), and testing (10%) sets.
Baseline Models. To comprehensively evaluate our model’s performance, we construct a series of strong baseline models by replacing key components of our proposed architecture. While ensuring a relatively fair comparison, we independently optimize the hyperparameters for each architecture. (i)Non-Graph Model.We replace the RGCN module with a Conditional GAN (CGAN) (Isola et al., 2017) to evaluate the utility of an explicit graph structure. (ii)Homogeneous Graph Models. We replace the RGCN module with GCN and GINE to assess the importance of modeling heterogeneous relationships. (iii)Generative Graph Models. We replace the RGCN with generative methods, including a Graph Autoencoder (GAE) and a Graph GAN (GGAN), where the CNN layers in a standard GAN are replaced with RGCN layers. (iv)Temporal Variants.We replace the LSTM module with a GRU and a Transformer encoder to analyze the impact of different temporal modeling components.
Evaluation Metrics. The predictive performance of the models is evaluated using three standard regression metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and the Coefficient of Determination (R²). To assess computational cost, we record the number of floating-point operations (FLOPs), training time (seconds per epoch), and inference speed (inferences per second). All models are trained using the Adam optimizer with a Mean Squared Error (MSE) loss function. For the GAE, the loss function is augmented with a KL divergence term; for the CGAN, a binary cross-entropy loss is added. To ensure robust training, we employ the ReduceLROnPlateau learning rate scheduling strategy and an early stopping mechanism.
@@ 6 Result
This section presents a comprehensive experimental evaluation of our proposed UrbanGraph framework. We begin in Section 6.1 by comparing its performance against various baseline models to validate its effectiveness. Subsequently, in Section 6.2, we conduct key ablation studies to quantify the contributions of our model’s core components. Further detailed analyses—including additional ablation studies, hyperparameter sensitivity, and a computational performance evaluation—are provided in Appendix E.
@@ 6.1 Model Performance
As shown in Table 1, our proposed UrbanGraph achieves the best performance across all evaluation metrics. On the test set, the model achieves the highest average R² of 0.8542 and the lowest RMSE of 1.0535, outperforming all baseline models. Compared to all baselines, it improves prediction accuracy by up to 10.8% (in R²) and enhances computational efficiency by 17.0% (in FLOPs). This highlights the effectiveness of the proposed method in efficiently capturing complex spatio-temporal dependencies.
@@
Model Flops Test Time Cost
Avg R² Avg RMSE Training (epoch/s) Inference/s
CGAN-LSTM $1.10\times 10^{10}$ $0.7712\pm.0369$ $1.3450\pm.1175$ $15.3252\pm 1.0999$ $1.5558\pm.1951$
GCN-LSTM $8.28\times 10^{9}$ $0.8347\pm.0039$ $1.1327\pm.0433$ $28.5321\pm 2.8358$ $2.8619\pm.4516$
GINE-LSTM $8.80\times 10^{9}$ $0.8087\pm.0226$ $1.2045\pm.0294$ $32.3169\pm 1.4643$ $3.1731\pm.2325$
GAE-LSTM $1.05\times 10^{10}$ $0.8494\pm.0036$ $1.0687\pm.0269$ $36.7376\pm 3.2079$ $3.6022\pm.4504$
GGAN-LSTM $9.44\times 10^{9}$ $0.8415\pm.0034$ $1.0981\pm.0406$ $42.4678\pm 3.1537$ $2.6488\pm.4073$
RGCN-GRU $7.12\times 10^{9}$ $0.8483\pm.0035$ $1.0682\pm.0380$ $20.8096\pm 1.3612$ $2.1640\pm.2133$
RGCN-Transformer $5.09\times 10^{10}$ $0.8465\pm.0065$ $1.0791\pm.0253$ $7.6463\pm.8325 3.3$$3345\pm.1482$
URBANGRAPH $9.13\times 10^{9}$ $\mathbf{0.8542\pm.0044}$ $\mathbf{1.0535\pm.0338}$ $\mathbf{24.4823\pm 0.9323}$ $\mathbf{2.6914\pm.1404}$
@@
In addition to these quantitative metrics, we provide qualitative visualizations of the predicted heat maps in Appendix F, which intuitively demonstrate the model’s ability to capture fine-grained spatial distributions.
Figure 4: Model performance analysis. (a) Training and validation R² convergence curves for the UrbanGraph model. Shaded areas represent the confidence interval. (b) Hour-by-hour R² comparison between UrbanGraph and baselines on the test set, with error bars indicating standard deviation.
@@ 6.2 Ablation Analysis
Heterogeneous Graph Mechanism. To validate the importance of modeling diverse physical interactions with distinct relation types, we compare our full model (Base), which uses a heterogeneous graph (RGCN), against a variant that simplifies the graph to be homogeneous (GCN). The results, shown in Table 2(a), reveal a significant performance degradation when heterogeneity is removed, with the R² score dropping from 0.8629 to 0.8347. This underscores the necessity of using a heterogeneous framework to differentiate between various physical processes.
Dynamic Graph Mechanism. To validate the effectiveness of the dynamic graph mechanism, we compare our model with a variant that uses a static graph (i.e., the same graph structure is shared across all timesteps). As shown in Table 2(b), disabling the dynamic mechanism leads to a significant performance drop in the model (Static), with the R² score decreasing from 0.8629 to 0.8057. This highlights that explicitly modeling the temporal evolution of spatial interactions is crucial for this prediction task.
@@ 7 Conclusion
In this paper, we proposed UrbanGraph, a physics-informed dynamic heterogeneous graph framework for solving urban dynamic heterogeneous graph computing tasks. We tested it on the problem of microclimate prediction, where UrbanGraph achieved the best performance compared to all baselines. It improves prediction accuracy by up to 10.8% (in R²) and computational efficiency by 17.0% (in FLOPs), with the heterogeneous and dynamic graph mechanisms contributing gains of 3.5% and 7.1%, respectively. Furthermore, the UMC4/12 dataset, which we constructed and released, serves as the first high-resolution benchmark in this field and will help accelerate the development and fair comparison of new algorithms in the future. In summary, our work advances the application of data-driven methods in the field of urban physical field prediction and points to a promising direction for future research.
Limitation and Future Work. Our work explicitly encodes predefined physical processes (i.e., prior knowledge) into the graph topology. While this has shown performance advantages, it may oversimplify the real physical processes, as it might overlook latent relationships present in the data that we have not yet modeled or are unknown. Therefore, a key direction for future research is to explore adaptive graph learning methods. The core objective is to design a framework that can automatically learn and optimize the graph structure from data in an end-to-end manner. Such a framework could not only surpass existing frameworks on prediction tasks but also uncover unknown interaction patterns among urban entities, providing new insights for urban science.
@@
To ensure reproducibility, our code and the UMC4/12 dataset are publicly available at https://github.com/wlxin-nus/UrbanGraph.
@@ Appendix A Key Physical Equations in ENVI-met
This appendix outlines the key physical equations within the ENVI-met model (Bruse & Fleer, 1998) used to generate the dataset for this study.
@@ Appendix B High-Resolution Spatio-Temporal Dataset for Microclimate and Thermal Comfort
We built the 3D model input files for the ENVI-met simulations by integrating multiple public geospatial data sources. Specifically, we resampled and performed 3D voxelization on building footprints and heights from OpenStreetMap (Haklay & Weber, 2008), a high-resolution land cover classification map (Gaw et al., 2019), and an ultra-high-resolution canopy height map (Tolan et al., 2024). This process generated ENVI-met input files (.INX) with a horizontal resolution of 4 meters and a vertical resolution of 3 meters. To ensure high fidelity, we assigned realistic material properties to different surfaces and building boundaries, and specified corresponding tree species for vegetation of varying heights. The detailed material assignments and parameters are provided in Appendix Table A2 and A3. The simulation period covered the hours from 08:00 to 19:00, when urban heat effects are most significant.
@@
Figure A1: Visualization of the input data for two sample sites from the UMC4/12 dataset. The top row shows a dense, mixed-use urban area, while the bottom row depicts a residential area with more green space. Each column represents a different data layer: (left) Tree Height, (middle) Land Cover Type, and (right) Building Height.
@@
Figure A2: Visualization of the spatio-temporal simulation output for a single urban block. Each row corresponds to a different target variable: AT, WS, RH, MRT, PET, and UTCI. Each column represents a specific hour, showing the dynamic evolution of the microclimate from morning (09:00) to evening (19:00).
@@ E.1 Detailed Ablation Studies
Table A7: Effectiveness of the Warming-up Mechanism.
Model MSE
Base 0.8629 1.0976
NP1 0.8510 1.1526

### v2

@@ 5 Experiments
Dataset. Experiments utilize a high-fidelity ENVI-met dataset where the environment is discretized into a 3D grid (44m horizontal, 33m vertical) derived from high-precision vector and land cover data (Appendix B.1). While primarily evaluating UTCI, we predict all six variables to demonstrate the architecture’s scalability. Temporal graphs follow principles in Section 4.1. The 396 urban blocks are split into Training (70%), Validation (20%), and Testing (10%). Crucially, the test set comprises spatially distinct, unseen blocks to strictly evaluate generalization to new configurations.
Baseline Models. We benchmark UrbanGraph against four state-of-the-art categories: 1) Grid-based Methods (CGAN-LSTM (Isola et al., 2017), Pix2Pix+PINN) to assess graph necessity, where the latter utilizes soft physics-loss constraints; 2) Static Spatio-Temporal GNNs (GCN/GINE-LSTM, STGCN (Yu et al., 2018), ASTGCN (Guo et al., 2019)) to validate dynamic topology against fixed structures; 3) Generative Graph Models (GGAN-LSTM, GAE-LSTM) to compare with latent structure learning; and 4) Dynamic Graph Models (LRGCN (Li et al., 2019b)) to demonstrate the efficiency of our explicit causal encoding over implicit recurrent learning.
Evaluation Metrics. We evaluate predictive performance using Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Coefficient of Determination (R²). Computational efficiency is assessed via floating-point operations (FLOPs), training time, and inference speed. Models are trained using Adam with Mean Squared Error (MSE) loss—augmented by a KL divergence term for GAE and binary cross-entropy for CGAN. Training employs ReduceLROnPlateau scheduling and early stopping for robustness.
@@ 6 Result
We evaluate UrbanGraph by strictly benchmarking it against baselines in Section 6.1 and conducting core ablation studies in Section 6.2. Extended analyses, including additional ablations, hyperparameter sensitivity, and computational efficiency, are detailed in Appendix E.
@@ 6.1 Model Performance
Figure 4: Input data configuration and prediction visualization. (a) Dynamic Global Features: Time-varying meteorological forcing (e.g., Wind Speed, Temperature, Relative Humidity). (b) Static GIS Features: Node-level spatial attributes including Building Height, Tree Height, and Land Cover ID. (c) Prediction Results: Visual comparison between Ground Truth, Model Prediction, and Absolute Error maps.
By integrating dynamic meteorological forcing with static GIS features, UrbanGraph achieves high-fidelity microclimate prediction across complex urban environments. As shown in Figure 4, visual comparisons confirm that the model accurately captures fine-grained thermal gradients, exhibiting strong spatial consistency with ground truth and minimal prediction error. To further substantiate these visual findings, we provide extended qualitative visualizations in Appendix F, which offer additional intuitive evidence of the model’s ability to capture complex spatial distributions.
Table 1 shows UrbanGraph achieves state-of-the-art performance (highest $R^{2}=0.8542$). Compared to the strongest dynamic graph baseline LRGCN ($R^{2}=0.8422$), UrbanGraph improves accuracy while reducing FLOPs by 73.8% ($9.13\times 10^{9}$ vs. $3.49\times 10^{10}$) and training time by 21% (24.5s vs. 31.1s), validating the efficiency of explicit causal pruning over implicit recurrence. Furthermore, surpassing Pix2Pix+PINN ($0.8320$) confirms that hard structural constraints offer better physical consistency than soft loss constraints, a finding corroborated by ablation studies in Appendix E.
@@
Category Model Flops Test Time Cost
Avg R² Avg RMSE Avg MAE Training (epoch/s) Inference/s
Grid- based CGAN-LSTM $1.10\times 10^{10}$ $0.7712\pm.0369$ $1.3450\pm.1175$ $0.9539\pm.0611$ $15.3252\pm 1.0999$ $1.5558\pm.1951$
Pix2Pix+PINN $1.10\times 10^{10}$ $0.8320\pm.0046$ $1.1485\pm.0306$ $0.8365\pm.0185$ $17.5020\pm 0.0576$ $1.3031\pm.0253$
Static STGNNs GCN-LSTM $8.28\times 10^{9}$ $0.8347\pm.0039$ $1.1327\pm.0433$ $0.8544\pm.0308$ $28.5321\pm 2.8358$ $2.8619\pm.4516$
GINE-LSTM $8.80\times 10^{9}$ $0.8087\pm.0226$ $1.2045\pm.0294$ $0.9030\pm.0190$ $32.3169\pm 1.4643$ $3.1731\pm.2325$
STGCN $\mathbf{4.13\times 10^{9}}$ $0.7880\pm.0065$ $1.2958\pm.0287$ $0.9879\pm.0243$ $\mathbf{5.3547\pm 0.0222}$ $\mathbf{0.2928\pm.0041}$
ASTGCN $2.06\times 10^{10}$ $0.8317\pm.0048$ $1.1491\pm.0234$ $0.8579\pm.0120$ $25.4938\pm 0.0438$ $1.2903\pm.0284$
Generative Graph GAE-LSTM $1.05\times 10^{10}$ $0.8494\pm.0036$ $1.0687\pm.0269$ $0.7968\pm.0128$ $36.7376\pm 3.2079$ $3.6022\pm.4504$
GGAN-LSTM $9.44\times 10^{9}$ $0.8415\pm.0034$ $1.0981\pm.0406$ $0.8214\pm.0319$ $42.4678\pm 3.1537$ $2.6488\pm.4073$
Dynamic Graph RGCN-GRU $7.12\times 10^{9}$ $0.8483\pm.0035$ $1.0682\pm.0380$ $0.8020\pm.0293$ $20.8096\pm 1.3612$ $2.1640\pm.2133$
RGCN-Transformer $5.09\times 10^{10}$ $0.8465\pm.0065$ $1.0791\pm.0253$ $0.8066\pm.0118$ $7.6463\pm.8325 3.3$$3345\pm.1482$
LRGCN $3.49\times 10^{10}$ $0.8422\pm.0061$ $1.0889\pm.0321$ $0.8100\pm.0160$ $31.0808\pm 0.1877$ $3.0550\pm.0348$
URBANGRAPH $9.13\times 10^{9}$ $\mathbf{0.8542\pm.0044}$ $\mathbf{1.0535\pm.0338}$ $\mathbf{0.7866\pm.0250}$ $24.4823\pm 0.9323$ $2.6914\pm.1404$
@@
Real-world generalization requires dynamic edge rules that can withstand observational noise. We verified this robustness through extensive multi-scale validation (Appendix A2). The model achieved high accuracy in both micro-scale calibration on the NUS campus ($r>0.73$) and city-scale deployment across Singapore ($r=0.842$).This successful generalization confirms that our parameterization strategy remains effective across heterogeneous urban morphologies.
To rigorously evaluate the model’s generalization across distinct physical domains, we constructed the UWF3D dataset, which records high-resolution vector fields governed by Navier-Stokes equations. UrbanGraph demonstrates exceptional adaptability to this new physics, achieving high accuracy ($R^{2}>0.88$ for the $u$-component) and significantly outperforming the Grid-GCN baseline. This successful transition from scalar thermal diffusion to complex vector flow dynamics confirms the robustness of our explicit topological encoding framework in capturing diverse physical mechanisms. Detailed setup and results are provided in Appendix G.
Figure 5: Model performance analysis. (a) Test set $R^{2}$ convergence curves for the UrbanGraph model. Shaded areas represent the confidence interval. (b) Hour-by-hour $R^{2}$ comparison between UrbanGraph and baselines on the test set, with error bars indicating standard deviation.
@@ 6.2 Ablation Analysis
Heterogeneous Graph Mechanism. To validate the importance of modeling diverse physical interactions with distinct relation types, we compare our full model (Base), which uses a heterogeneous graph (RGCN), against a variant that simplifies the graph to be homogeneous (GCN). The results, shown in Table 2(a), reveal a significant performance degradation when heterogeneity is removed, with the R² score dropping from 0.8629 to 0.8347. The performance drop highlights causal entanglement in homogeneous graphs where a single parameter set fits distinct physical laws. Heterogeneity resolves this by enabling physical operator decoupling.
Dynamic Graph Mechanism. To validate the effectiveness of the dynamic graph mechanism, we compare our model with a variant that uses a static graph (i.e., the same graph structure is shared across all timesteps). As shown in Table 2(b), disabling the dynamic mechanism leads to a significant performance drop in the model (Static), with the R² score decreasing from 0.8629 to 0.8057. The gain validates time-varying causal pruning: unlike static graphs, the dynamic mechanism actively removes physically irrelevant connections (e.g., shifting shadows), thereby reducing optimization difficulty.
@@ 7 Conclusion
In this paper, we proposed UrbanGraph, a physics-informed dynamic graph framework for microclimate prediction. Achieving state-of-the-art accuracy ($R^{2}=0.8542$), UrbanGraph outperforms the strongest baseline (LRGCN) with a 73.8% reduction in FLOPs and 21% faster training. This result validates the superior efficiency of explicit causal pruning over implicit recurrence.
Furthermore, the UMC4/12 dataset, which we constructed and released, serves as the first high-resolution benchmark in this field and will help accelerate the development and fair comparison of new algorithms in the future. In summary, UrbanGraph demonstrates the potential of structural priors in bridging the gap between numerical simulation and deep learning. Beyond microclimate prediction, our work offers a generalizable explicit topological encoding paradigm for spatio-temporal dynamics governed by known physical equations.
Limitation and Future Work. Our work explicitly encodes predefined physical processes (i.e., prior knowledge) into the graph topology. While this has shown performance advantages, it may oversimplify the real physical processes, as it might overlook latent relationships present in the data that we have not yet modeled or are unknown. To address this trade-off between physical consistency and model flexibility, a critical future direction is to develop a Hybrid Topology framework. By augmenting the fixed physical graph with a sparse, learnable Residual Graph, future models could uncover unmodeled interaction patterns from data residuals. This approach effectively balances the robustness of engineering rules with the flexibility required to identify new physical mechanisms.
@@
To ensure reproducibility, our code and the UMC4/12 dataset are publicly available at https://github.com/wlxin-nus/UrbanGraph.git.
@@ Appendix A Key Physical Equations in ENVI-met
This appendix outlines the key physical equations within the ENVI-met model (Bruse and Fleer, 1998) used to generate the dataset for this study.
@@ B.1 Dataset Generation
We built the 3D model input files for the ENVI-met simulations by integrating multiple public geospatial data sources. Specifically, we resampled and performed 3D voxelization on raw data with varying precisions: building footprints were extracted from vector-based OpenStreetMap data (Haklay and Weber, 2008), while the land cover classification (Gaw et al., 2019) and canopy height maps (Tolan et al., 2024) utilized a 1m high resolution. This process generated ENVI-met input files (.INX) with a uniform horizontal resolution of 4 meters and a vertical resolution of 3 meters. To ensure high fidelity, we assigned realistic material properties to different surfaces and building boundaries, and specified corresponding tree species for vegetation of varying heights. The detailed material assignments and parameters are provided in Appendix Table A2 and A3. The simulation period covered the hours from 08:00 to 19:00, when urban heat effects are most significant.
@@
Figure A1: Visualization of the input data for ten sample sites from the UMC4/12 dataset. For every site, three data layers are visualized: (left in the triplet) Tree Height, (middle) Land Cover Type, and (right) Building Height.
@@
B.2 REAL-WORLD VALIDATION STRATEGY
To ensure that our physics-informed framework generalizes to real-world scenarios, we implemented a two-tier validation process covering both micro-scale fidelity and city-scale consistency.
B.2.1 Micro-scale Calibration with Field Measurements
We conducted a rigorous validation of the ENVI-met model based on a sensor network deployed on the campus of the National University of Singapore (NUS).
Figure A2: Visualization of the spatio-temporal simulation output for a single urban block. Each row corresponds to a different target variable: AT, WS, RH, MRT, PET, and UTCI. Each column represents a specific hour, showing the dynamic evolution of the microclimate from morning (09:$0$) to evening (19:$0$).
Figure A3: ENVI-met model validation. (a) Map of the validation area and sensor locations on the NUS campus. (b) Comparison of simulated vs. measured diurnal cycles for Temperature, Relative Humidity, and Wind Speed, showing strong agreement ($r>0.73$).
As shown in Figure A3a, a sensor on an open rooftop (Reference Point) was selected to provide the driving meteorological inputs. We compared simulation results with measured data from five distributed sensor stations covering diverse morphologies.
The plots in Figure A3b show the hourly variations of key variables: air temperature, relative humidity, and gust wind speed. The results indicate high agreement with measured values (Pearson’s correlation coefficient $r>0.73$, $p<0.01$), accurately capturing diurnal trends. This confirms that our simulation setup reliably reproduces key microclimate dynamics in complex urban environments, providing a solid physical basis for subsequent data-driven modeling.
B.2.2 City-scale Cross-Validation
To further assess the model’s dynamic response capability in the temporal dimension and its generalization at the city scale, we introduced the HiGTS dataset (Jian et al., 2024), which provides global hourly UTCI at a $0.1^{\circ}\times 0.1^{\circ}$ resolution.
Figure A4: City-scale generalization validation against the HiGTS dataset. The scatter plot compares UrbanGraph predictions with HiGTS reference hourly UTCI for Singapore. The results demonstrate a strong correlation ($r=0.842$) and low bias, confirming the model’s robustness in generalizing to real-world urban environments.
We deployed the trained UrbanGraph model to predict the hourly UTCI for the entire Singapore island on May 3rd, 2023, and compared the aggregated results with the HiGTS reference values.
As shown in the scatter plot in Figure A4, the two datasets show a strong correlation ($r=0.842$), with a slight systematic underestimation (Bias = $0.667^{\circ}C$). The hourly analysis reveals that prediction bias primarily originates from trend discrepancies between the simulation inputs and the reference dataset at specific moments (e.g., abrupt weather changes) rather than flaws in the model itself. This validates that UrbanGraph can effectively generalize to real-world urban data beyond the training distribution.
@@ C.1 Parameter Rationale
The boundary parameters for these edges are structural constraints, with their values set by referencing classical urban climatology studies (Ziter et al., 2019; Dare, 2005; Ng, 2009; Lin et al., 2023). This setting is necessary for explicit physical embedding, as it translates domain knowledge (e.g., the effective decay distance of vegetation cooling) into reliable physical upper bounds for most typical urban environments. Within these bounds, the actual interaction strength is adaptively learned by the network. The robustness of this setup is empirically validated by its successful generalization to the UWF3D dataset (Appendix G), where these physical bounds yielded high accuracy on a completely different physical task.
@@ E.1 Detailed Ablation Studies
Physics Injection Strategy. To rigorously isolate the impact of different physics-injection paradigms, we conducted a controlled experiment. We compared three variants sharing the identical backbone architecture (GCN+LSTM) and hyperparameter settings (e.g., hidden dimensions, learning rate). The only difference lies in the mechanism of physics injection: 1) Baseline: A standard data-driven model without any physical constraints. 2) Soft Constraint: The baseline augmented with PDE-based loss functions (PINN approach). 3) Hard Constraint (Ours): The proposed UrbanGraph that encodes physics directly into the graph topology.
Table A6: Quantitative comparison of different physics injection strategies.
Strategy Flops Test Time Cost
Avg R² Avg RMSE Avg MAE Training (epoch/s) Inference/s
Baseline (Data-driven) $8.28\times 10^{9}$ $0.7846\pm.0039$ $1.2918\pm.0389$ $0.9953\pm.0248$ $\mathbf{11.9636\pm 0.0046}$ $\mathbf{1.0002\pm.0082}$
Soft Constraint (PINN) $8.28\times 10^{9}$ $0.7876\pm.0023$ $1.3028\pm.0531$ $1.0037\pm.0359$ $12.0993\pm 0.0127$ $1.0164\pm.0123$
Hard Constraint (Ours) $9.13\times 10^{9}$ $\mathbf{0.8542\pm.0044}$ $\mathbf{1.0535\pm.0338}$ $\mathbf{0.7866\pm.0250}$ $24.4823\pm 0.9323$ $2.6914\pm.1404$
As shown in Table A6, the ’Soft Constraint’ yields negligible improvement ($\Delta R^{2}\approx+0.0030$) over the baseline. In contrast, the ’Hard Constraint’ achieves a substantial performance leap ($\Delta R^{2}\approx+0.0696$). Although constructing the dynamic topology incurs a computational cost (increasing training time from  12s to 24.5s), the return on investment is significant: The hard structural constraint delivers over 23 times the accuracy gain of the soft loss constraint.
@@
Table A9: Effectiveness of the Warming-up Mechanism.
Model MSE
Base 0.8629 1.0976
NP1 0.8510 1.1526